# On the magnetic field evolution of interplanetary coronal mass ejections from 0.07 to 5.4 au

### v1

@@ Abstract
A central question for understanding interplanetary coronal mass ejection (ICME) physics and improving space weather forecasting is how ICMEs evolve in interplanetary space. We have updated one of the most comprehensive in situ ICME catalogs to date, which now includes 1976 events from 11 space missions covering over 34 years, from December 1990 to August 2025. We have combined existing catalogs including magnetic obstacles and identified and added boundaries of an additional 807 (40.8%) events ourselves. With this catalog, we demonstrate the most extensive analysis to date of total ICME magnetic field values as a function of heliocentric distance. Parker Solar Probe has observed 6 ICMEs at <0.23<0.23 au (until April 2025), and Solar Orbiter and BepiColombo have added more events near 0.3 au, bridging the major observational gap towards the solar corona. Our main result is that a single power law can describe the evolution of the mean total magnetic field (exponent value of k=1.57k=-1.57) and maximum field (k=1.53k=-1.53) in the magnetic obstacle (MO), from 0.07 to 5.4 au. Extending the power law to the solar photosphere reveals a strong inconsistency with magnetic field magnitudes observed in the quiet Sun and active regions by 2 and 4 orders of magnitude, respectively. We introduce a multipole-type power law with two exponents, $k_{1}=-1.57$, and $k_{2}=-6$, relating the ICME magnetic field magnitude to an average solar active region field strength. These results present important observational constraints for the evolution of ICMEs from the Sun to the heliosphere.
@@
1 Introduction
@@
Decades of observations of interplanetary coronal mass ejections (ICMEs) have still left us with many unsolved puzzles. Some of the key ones among them are the 3D magnetic structure of the flux rope and its extent in all three dimensions. Furthermore, how they evolve after their launch from the Sun towards interplanetary space, passing the planets and moving towards the outer heliosphere is still not fully understood (e.g. manchester2017; luhmann_2020_review; temmer2023; alhaddad2025structure). With the Parker Solar Probe (PSP; Fox_2016) mission, launched in 2018, a new region below 0.3 au has been unlocked for the first time for in situ observations of ICMEs (e.g. nieves_chinchilla_2020; winslow_2021_sta_psp; salman_2024_psp_cat; davies2024; trotta2024; jebaraj2024).
@@
Here, we combine these in situ observations of ICMEs by PSP with data from ten other spacecraft missions to form a catalog that is as uniform as possible. Our ICMECAT catalog contains 1976 ICMEs observed in situ, covers 3 solar cycles from 1990 to 2025, and heliocentric distances from 0.07–5.4 au. This means we present the most comprehensive analysis to date of the radial evolution of the total magnetic field strength of ICMEs measured in situ. We put a particular emphasis on their magnetic obstacle intervals, which often contain magnetic flux ropes (e.g. cane2003) that drive strong geomagnetic storms with their potential long-duration southward magnetic fields (e.g. zhang2007). This information is of major importance for developing and validating empirical and numerical models of ICME flux ropes for many heliophysics applications.
@@
Many authors have considered power law fits for the relationship between the total magnetic field in ICMEs against distance from the Sun (e.g., leitner_2007; gulisano2007; good2016; Hanneson2020; davies2021_juno; davies2022; salman_2024_psp_cat). Statistical results indicate power laws with exponents $k$ in the range between k=1k=-1 and k=2k=-2, depending on the heliocentric distance range covered. However, studies that use multi-spacecraft conjunctions for the same event at different heliocentric distances paint a picture where individual events can differ from the average parameters of the power laws (e.g. leitner_2007; Salman_2020; davies2022). For example, Davies_2021_solo demonstrate a case of almost no radial expansion between 0.8 and 1.0 au, as the slow CME that they study is pushed by a high-speed stream from behind and thus constrained its expansion. Even with individual events deviating from these empirical relationships, the laws are a highly useful basis to describe the physics of CME evolution with empirical models.
@@
Our intent with this study is not only to improve the understanding of the physics of CMEs, but the results are highly useful for practical space weather applications, as there is now renewed interest in the concept of so-called upstream or sub-L1 monitors (see lindsay1999; cyr2000diamond; kubicka2016; laker2024; lugaz2025need; weiler2025; davies2025_solo). These are spacecraft that measure the magnetic field strength at a position closer to the Sun than the Sun–Earth Lagrange 1 (L1) point, but still near the Sun–Earth line, providing earlier measurements of the ICME magnetic field structure than current space weather monitors at L1, extending the forecast lead time. The European Space Agency (ESA) is currently building the HENON spacecraft (e.g. cicalo2025henon), planned to be launched at the end of 2026. This mission acts as a pathfinder for ESA’s proposed sub-L1 mission concept, SHIELD, which plans to deploy multiple spacecraft on a distant retrograde orbit (henon1969; perozzi2017) at heliocentric distances of about 0.85–0.9 au. Power-law scaling of ICME magnetic fields provides a straightforward method to extrapolate the observations of these sub-L1 spacecraft to Earth distance.
@@
Power-law descriptions of ICME magnetic fields are also crucial for Farady rotation methods, which use radio observations to remotely probe the magnetic structure of ICMEs (e.g. wood2020faraday; kooi2022_faraday; jensen2025polarized). These methods require assumptions about how the magnetic field decreases with heliocentric distance. Therefore, accurate power laws are essential for converting observed Faraday rotation into predictions of magnetic field strengths at larger heliocentric distances. This is particularly important as Faraday rotation offers one of the few options to obtain information about a CMEs internal magnetic field structure before it is measured in situ.
Our ICMECAT catalog (Moestl2017; moestl2020) has grown over the recent years into a unified and comprehensive database of in situ ICMEs observed by 11 spacecraft over more than 3 solar cycles. It has been widely used in different contexts, not only for case studies, but also for studying intermittency in CMEs (ruohotie2025), or training machine learning algorithms for automatic identification of ICMEs in solar wind in situ observations (nguyen2019; ruedisser2022; ruedisser2025arcane). Its main purpose is to identify multi-spacecraft observations of CMEs to advance our knowledge of their 3D evolution, shape and structure (e.g. moestl2022; regnault2024; palmerio2025event4; alhaddad2025structure; zhang2025_multipoint).
2 Spacecraft observations
@@
2.1 ICME catalog
At the heart of our analysis is the ICMECAT, an open source and living ICME catalog which is accessible online (see Acknowledgments section), presented so far in Moestl2017 and moestl2020. A search for multipoint ICME events using the ICMECAT has been demonstrated in moestl2022. Here, we present a major update that now acts as the main reference publication for the ICMECAT, and briefly summarize its main features. We have significantly extended the catalog in comparison to previous work, with the main goal of adding more ICME events observed in the inner heliosphere (<1.0<1.0 au) that have recently become available, possible due to the coverage of more recent spacecraft missions. Our aim is to provide a living database to better understand CME magnetic field evolution, which rapidly includes novel events soon after they are observed. With respect to the solar activity in solar cycle 25, rising from solar minimum around the launches of PSP in 2018 and Solar Orbiter in 2020 to solar maximum around 2024, the updated catalog now includes ICMEs for about half of solar cycle 25. It also covers the first 23 orbits of PSP until 2025 April 30, of which two orbits reached the planned closest approach distance of the mission (0.0458 au).
Figure 1a shows the heliocentric distance at which the ICMECAT events were observed over time, where the different color markers correspond to the observing spacecraft. The ICMECAT now covers 1976 ICME events from the years 1990 to 2025. Figure 1b gives an overview of the longitudinal distribution of events in Heliocentric Earth Equatorial (HEEQ) coordinates against radial distance, where the Earth is fixed on the right side at 0° longitude. This demonstrates how many ICMEs have been observed in the space age after 1990, in particular around 1 au by the STEREO mission and through several missions venturing into the inner heliosphere, including both planetary spacecraft (e.g. MESSENGER, Venus Express, and BepiColombo) and dedicated solar wind missions (e.g. PSP and Solar Orbiter). Figures 1c and 1d show a subset of events with the (predicted) trajectories of PSP and Solar Orbiter, respectively, as well as the more recent ICME events, with PSP observing events at distances <0.3<0.3 au at almost every recent close approach to the Sun. BepiColombo is measuring ICMEs with magnetic field instrument MPO-MAG (heyner2021) until orbit insertion at Mercury in 2026, and Solar Orbiter is now starting to provide high latitude ICME observations since raising its orbital inclination after February 2025.
@@ 2.1 ICME catalog
nieves_chinchilla_2020 1995 February – 2015 December 342 Wind
@@
jian2018 2007 January – 2017 December 188 STEREO-A
jian2018 2007 January – 2014 September 150 STEREO-B
davies2021_juno; davies2022 2011 September – 2016 May 53 Juno
@@
good2018 2007 February – 2013 December 80 VEX
@@
good2018 2007 May – 2011 November 22 MESSENGER
winslow2018 2011 May – 2014 September 55 MESSENGER
@@
richardson2014identification 1990 December – 2007 September 279 Ulysses
@@
The ICMECAT is an aggregate of catalogs published by nieves_chinchilla_2020 for Wind, jian2018 for STEREO-A/B, davies2021_juno; davies2022 for Juno, good2018 for Venus Express and MESSENGER, winslow2018 for MESSENGER and richardson2014identification for Ulysses. In addition, we have added ICME events identified by ourselves for the missions STEREO-A, Wind, PSP, Solar Orbiter, BepiColombo and a few events for MAVEN, hence including altogether 11 different spacecraft missions in the ICMECAT. We note that we have also adjusted some event times originally taken from the existing STEREO-A and Wind mission catalogs. In total, 807 events (40.8% of the catalog) were included in addition to existing catalog events or modified by our team. In Table 1 we summarize the catalog sources.
@@
When combining different catalogs, one needs to be careful with the criteria that were used to identify ICME events. While the catalogs that we use do not match perfectly in their criteria, all catalogs that we have used include magnetic obstacles (MOs, also known as magnetic ejecta). Shock arrivals, followed by sheath regions but lacking MOs, are not included in the ICMECAT, however, events with only MOs and no sheaths are included. All events were selected visually according to the criteria given below, namely for both the events given in the existing catalogs and those identified by ourselves. Automatic identification methods (e.g. camporeale2017; nguyen2019; ruedisser2022; nguyen2025_multiclass; ruedisser2025arcane) are now reaching a promising level, but for consistency, we stick to manual identification for the moment.
@@
All MOs in the ICMEs are broadly characterized by a high total magnetic field strength and a smooth rotation of the magnetic field vector. If plasma data is present, we look for additional signatures such as declining speed profiles and possibly low proton temperature within the MO, which are signatures associated with expansion of the MO. Each event has three times assigned to it: (1) ICME start time, (2) MO start time, and (3) MO end time. For interacting ICMEs, we try to identify individual MOs as standalone events (e.g. for the May 2024 ICMEs, see weiler2025). If no shock is present, in our catalog the ICME start time reflects the beginning of a significant density enhancement in front of the MO. If neither is present, the ICME start time is set equal to the MO start time. It is well known that different human observers often disagree on the exact times of the MO interval and the ICME start time (if no shock is present). Nevertheless, these uncertainties that are inherently present in these types of catalogs will have only little influence on the derivation of the global statistical results that we present here, although we note that distributions of general parameters at 1 au can be affected (see kay2025collection).
@@
Concerning the in situ data, we use the magnetic field components and total field, as well as bulk proton data (speed, temperature, density, if available), linearly interpolated to a 1 minute resolution, in order to calculate parameters and plot a figure for each event in a computationally efficient way. For missions where the available cadence is greater than 1 minute, we use a different time resolution. For Wind, we interpolate the data to 2 minutes. For Ulysses, an 1 hour combined plasma and magnetic field dataset is readily available. For the few MAVEN events that we include, we interpolate to one data point per orbit. This results in a largely homogeneous dataset for all 11 spacecraft with a total size of only about 10 gigabyte. The catalog is then created by reading in all available spacecraft data, and another file containing the heliospheric positions of each spacecraft produced using the SPICE kernels available from most missions (see figshare links in the data section at the end of this study). We use the spacecraft position files to provide an overview of the distribution of the various spacecraft in the inner heliosphere at the time each of the ICME events is observed (see the ICMECAT figure database on the website and on figshare). We then calculate 37 parameters for each ICME based only on the 3 boundary times defined for each event (ICME start, MO start and MO end). In this way, all parameters are fully consistently recalculated every time the catalog is updated. We do not rely on any parameters from any previously published catalog, except for the 3 times that define an event, and calculate all cataloged ICME parameters by ourselves.
2.2 An example ICME observed by Solar Orbiter

### v2

# arXiv:2512.04730v2[astro-ph.SR] 02 Mar 2026

@@ Abstract
A central question for understanding interplanetary coronal mass ejection (ICME) physics and improving space weather forecasting is how ICMEs evolve in interplanetary space. We have updated one of the most comprehensive in situ ICME catalogs to date, which now includes 1976 events from 11 space missions covering over 34 years, from December 1990 to August 2025. We have combined existing catalogs including magnetic obstacles and identified and added boundaries of an additional 807 (40.8%) events ourselves. With this catalog, we demonstrate the most extensive analysis to date of total ICME magnetic field values as a function of heliocentric distance. Parker Solar Probe has observed 6 ICMEs at <0.23<0.23 au (until April 2025), and Solar Orbiter and BepiColombo have added more events near 0.3 au, bridging the major observational gap towards the solar corona. Our main result is that a single power law can describe the evolution of the mean total magnetic field (exponent value of k=1.57k=-1.57) and maximum field (k=1.53k=-1.53) for ICMEs with magnetic obstacles (MOs), from 0.07 to 5.4 au. Extending the power law to the solar photosphere reveals a strong inconsistency with magnetic field magnitudes observed in the quiet Sun and active regions by 2 and 4 orders of magnitude, respectively. We introduce a multipole-type power law with two exponents, $k_{1}=-1.57$, and $k_{2}=-6$, relating the ICME magnetic field magnitude to an average solar active region field strength. These results present important observational constraints for the evolution of ICMEs from the Sun to the heliosphere.
@@
software: numpy (Harris et al., 2020), pandas (McKinney, 2010; The Pandas Development Team, 2020), spacepy (Niehof et al., 2022), astropy (Astropy Collaboration et al., 2013, 2018, 2022), spiceypy (Annex et al., 2020), scipy (Virtanen et al., 2020), plotly (Kruchten et al., 2026) and matplotlib (Hunter, 2007).
I Introduction
@@
Decades of observations of interplanetary coronal mass ejections (ICMEs) have still left us with many unsolved puzzles. Some of the key ones among them are the 3D magnetic structure of the flux rope and its extent in all three dimensions. Furthermore, how they evolve after their launch from the Sun towards interplanetary space, passing the planets and moving towards the outer heliosphere is still not fully understood (e.g. Manchester et al., 2017; Luhmann et al., 2020; Temmer et al., 2023; Al-Haddad and Lugaz, 2025). With the Parker Solar Probe (PSP; Fox et al., 2016) mission, launched in 2018, a new region below 0.3 au has been unlocked for the first time for in situ observations of ICMEs (e.g. Nieves-Chinchilla et al., 2020; Winslow et al., 2021; Salman et al., 2024; Davies et al., 2024; Trotta et al., 2024; Jebaraj et al., 2024).
@@
Here, we combine these in situ observations of ICMEs by PSP with data from ten other spacecraft missions to form a catalog that is as uniform as possible. Our ICMECAT catalog contains 1976 ICMEs observed in situ, covers 3 solar cycles from 1990 to 2025, and heliocentric distances from 0.07–5.4 au. This means we present the most comprehensive analysis to date of the radial evolution of the total magnetic field strength of ICMEs measured in situ. We put a particular emphasis on their magnetic obstacle intervals, which often contain magnetic flux ropes (e.g. Cane and Richardson, 2003) that drive strong geomagnetic storms with their potential long-duration southward magnetic fields (e.g. Zhang et al., 2007). This information is of major importance for developing and validating empirical and numerical models of ICME flux ropes for many heliophysics applications.
@@
Many authors have considered power law fits for the relationship between the total magnetic field in ICMEs against distance from the Sun (e.g., Leitner et al., 2007; Gulisano et al., 2007; Good and Forsyth, 2016; Hanneson et al., 2020; Davies et al., 2021b, 2022; Salman et al., 2024). Statistical results indicate power laws with exponents $k$ in the range between k=1k=-1 and k=2k=-2, depending on the heliocentric distance range covered. However, studies that use multi-spacecraft conjunctions for the same event at different heliocentric distances paint a picture where individual events can differ from the average parameters of the power laws (e.g. Leitner et al., 2007; Salman et al., 2020; Davies et al., 2022). For example, Davies et al. (2021a) demonstrate a case of almost no radial expansion between 0.8 and 1.0 au, as the slow CME that they study is pushed by a high-speed stream from behind and thus constrained its expansion. Even with individual events deviating from these empirical relationships, the laws are a highly useful basis to describe the physics of CME evolution with empirical models.
@@
Our intent with this study is not only to improve the understanding of the physics of CMEs, but the results are highly useful for practical space weather applications, as there is now renewed interest in the concept of so-called upstream or sub-L1 monitors (see Lindsay et al., 1999; St. Cyr et al., 2000; Kubicka et al., 2016; Laker et al., 2024; Lugaz et al., 2025; Weiler et al., 2025; Davies et al., 2026). These are spacecraft that measure the magnetic field strength at a position closer to the Sun than the Sun–Earth Lagrange 1 (L1) point, but still near the Sun–Earth line, providing earlier measurements of the ICME magnetic field structure than current space weather monitors at L1, extending the forecast lead time. The European Space Agency (ESA) is currently building the HENON spacecraft (e.g. Cicalò et al., 2025), planned to be launched at the end of 2026. This mission acts as a pathfinder for ESA’s proposed sub-L1 mission concept, SHIELD, which plans to deploy multiple spacecraft on a distant retrograde orbit (Henon, 1969; Perozzi et al., 2017) at heliocentric distances of about 0.85–0.9 au. Power-law scaling of ICME magnetic fields provides a straightforward method to extrapolate the observations of these sub-L1 spacecraft to Earth distance.
@@
Power-law descriptions of ICME magnetic fields are also crucial for Farady rotation methods, which use radio observations to remotely probe the magnetic structure of ICMEs (e.g. Wood et al., 2020; Kooi et al., 2022; Jensen et al., 2025). These methods require assumptions about how the magnetic field decreases with heliocentric distance. Therefore, accurate power laws are essential for converting observed Faraday rotation into predictions of magnetic field strengths at larger heliocentric distances. This is particularly important as Faraday rotation offers one of the few options to obtain information about a CMEs internal magnetic field structure before it is measured in situ.
Our ICMECAT catalog (Möstl et al., 2017, 2020) has grown over the recent years into a unified and comprehensive database of in situ ICMEs observed by 11 spacecraft over more than 3 solar cycles. It has been widely used in different contexts, not only for case studies, but also for studying intermittency in CMEs (Ruohotie et al., 2025), for training machine learning algorithms for automatic identification of ICMEs in solar wind in situ observations (Nguyen et al., 2019; Rüdisser et al., 2022, 2026), or for removing ICMEs to study the radial evolution of solar wind parameters (Yogesh et al., 2026). Its main purpose is to identify multi-spacecraft observations of CMEs to advance our knowledge of their 3D evolution, shape and structure (e.g. Möstl et al., 2022; Regnault et al., 2024; Davies et al., 2024; Palmerio et al., 2025; Al-Haddad and Lugaz, 2025; Zhang et al., 2025).
II Spacecraft observations
@@
II.1 ICME catalog
At the heart of our analysis is the ICMECAT, an open source and living ICME catalog which is accessible online (see Data Availability Statement), presented so far in Möstl et al. (2017) and Möstl et al. (2020). A search for multipoint ICME events using the ICMECAT has been demonstrated in Möstl et al. (2022). Here, we present a major update that now acts as the main reference publication for the ICMECAT, and briefly summarize its main features. We have significantly extended the catalog in comparison to previous work, with the main goal of adding more ICME events observed in the inner heliosphere (<1<1 au) that have recently become available, possible due to the coverage of more recent spacecraft missions. Our aim is to provide a living database to better understand CME magnetic field evolution, which rapidly includes novel events soon after they are observed. With respect to the solar activity in solar cycle 25, rising from solar minimum around the launches of PSP in 2018 and Solar Orbiter in 2020 to solar maximum around 2024, the updated catalog now includes ICMEs for about half of solar cycle 25. It also covers the first 23 orbits of PSP until 2025 April 30, of which two orbits reached the planned closest approach distance of the mission (0.0458 au).
Figure 1a shows the heliocentric distance at which the ICMECAT events were observed over time, where the different color markers correspond to the observing spacecraft. The ICMECAT now covers 1976 ICME events from the years 1990 to 2025. Figure 1b gives an overview of the longitudinal distribution of events in Heliocentric Earth Equatorial (HEEQ) coordinates against radial distance, where the Earth is fixed on the right side at 0° longitude. This demonstrates how many ICMEs have been observed in the space age after 1990, in particular around 1 au by the STEREO mission and through several missions venturing into the inner heliosphere, including both planetary spacecraft (e.g. MESSENGER, Venus Express, and BepiColombo) and dedicated solar wind missions (e.g. PSP and Solar Orbiter). Figures 1c and 1d show a subset of events with the (predicted) trajectories of PSP and Solar Orbiter, respectively, as well as the more recent ICME events, with PSP observing events at distances <0.3<0.3 au at almost every recent close approach to the Sun. BepiColombo is measuring ICMEs with magnetic field instrument MPO-MAG (Heyner et al., 2021) until orbit insertion at Mercury in 2026, and Solar Orbiter is now starting to provide high latitude ICME observations since raising its orbital inclination after February 2025.
@@ II.1 ICME catalog
Nieves-Chinchilla et al. (2020) 1995 February – 2015 December 342 Wind
@@
Jian et al. (2018) 2007 January – 2017 December 188 STEREO-A
Jian et al. (2018) 2007 January – 2014 September 150 STEREO-B
Davies et al. (2021b, 2022) 2011 September – 2016 May 53 Juno
@@
Good et al. (2018) 2007 February – 2013 December 80 VEX
@@
Good et al. (2018) 2007 May – 2011 November 22 MESSENGER
Winslow et al. (2018) 2011 May – 2014 September 55 MESSENGER
@@
Richardson (2014) 1990 December – 2007 September 279 Ulysses
@@
The ICMECAT is an aggregate of catalogs published by Nieves-Chinchilla et al. (2020) for Wind, Jian et al. (2018) for STEREO-A/B, Davies et al. (2021b, 2022) for Juno, Good et al. (2018) for Venus Express and MESSENGER, Winslow et al. (2018) for MESSENGER and Richardson (2014) for Ulysses. In addition, we have added ICME events identified by ourselves for the missions STEREO-A, Wind, PSP, Solar Orbiter, BepiColombo and a few events for MAVEN, hence including altogether 11 different spacecraft missions in the ICMECAT. We note that we have also adjusted some event times originally taken from the existing STEREO-A and Wind mission catalogs. In total, 807 events (40.8% of the catalog) were included in addition to existing catalog events or modified by our team. In Table 1 we summarize the catalog sources.
@@
When combining different catalogs, one needs to be careful with the criteria that were used to identify ICME events. While the catalogs that we use do not match perfectly in their criteria, all catalogs that we have used include magnetic obstacles (MOs, also known as magnetic ejecta). Shock arrivals, followed by sheath regions but lacking MOs, are not included in the ICMECAT, however, events with only MOs and no sheaths are included. All events were selected visually according to the criteria given below, namely for both the events given in the existing catalogs and those identified by ourselves. Automatic identification methods (e.g. Camporeale et al., 2017; Nguyen et al., 2019; Rüdisser et al., 2022; Nguyen et al., 2025; Rüdisser et al., 2026) are now reaching a promising level, but for consistency, we stick to manual identification for the moment.
@@
All MOs in the ICMEs are broadly characterized by a high total magnetic field strength and a smooth rotation of the magnetic field vector. If plasma data is present, we look for additional signatures such as declining speed profiles and possibly low proton temperature within the MO, which are signatures associated with expansion of the MO. Each event has three times assigned to it: (1) ICME start time, (2) MO start time, and (3) MO end time. For interacting ICMEs, we try to identify individual MOs as standalone events (e.g. for the May 2024 ICMEs, see Weiler et al., 2025). A summary of the identification criteria is given in Table 2.
Table 2: Summarized criteria for interplanetary coronal mass ejection (ICME) identification.
Boundary Observable Signature Certainty
ICME start time Shock (sudden jumps in magnetic field, speed, density, temperature) high (shocks)
or start of density enhancement; if none, similar to MO start time average (density only)
MO start time Magnetic field data: start of elevated field strength, smooth rotation of low, transition of sheath to
field vectors over several hours, low field variance; look for discontinuities flux rope often unclear
Plasma data: Start of low proton temperature, low proton density,
linear declining speed profile
MO end time Magnetic field data: end of elevated field strength, smooth rotation of average, certainty of
field vectors over several hours, low field variance; look for discontinuities boundary for transition to
Plasma data: End of low proton temperature, low proton density, ambient solar wind varies
linear declining speed profile
If no shock is present, in our catalog the ICME start time reflects the beginning of a significant density enhancement in front of the MO. If neither is present, the ICME start time is set equal to the MO start time. It is well known that different human observers often disagree on the exact times of the MO interval and the ICME start time (if no shock is present). Nevertheless, these uncertainties that are inherently present in these types of catalogs will have only little influence on the derivation of the global statistical results that we present here, although we note that distributions of general parameters at 1 au can be affected (see Kay et al., 2026).
@@
Concerning the in situ data, we use the magnetic field components and total field, as well as bulk proton data (speed, temperature, density, if available), linearly interpolated to a 1 minute resolution, in order to calculate parameters and plot a figure for each event in a computationally efficient way. For missions where the available cadence is greater than 1 minute, we use a different time resolution. For Wind, we interpolate the data to 2 minutes. For Ulysses, an 1 hour combined plasma and magnetic field dataset is readily available. For the few MAVEN events that we include, we interpolate to one data point per orbit. This results in a largely homogeneous dataset for all 11 spacecraft with a total size of only about 10 gigabyte. The catalog is then created by reading in all available spacecraft data, and another file containing the heliospheric positions of each spacecraft produced using the SPICE kernels available from most missions (see figshare links in the data section at the end of this study). We use the spacecraft position files to provide a figure with an overview of the distribution of the various spacecraft in the inner heliosphere at the time each of the ICME events is observed (see the ICMECAT figure database on the website and on figshare). We then calculate up to 37 parameters for each ICME based only on the 3 boundary times defined for each event (ICME start, MO start and MO end). In this way, all parameters are fully consistently recalculated every time the catalog is updated. We do not rely on any parameters from any previously published catalog, except for the 3 times that define an event, and calculate all cataloged ICME parameters by ourselves.
II.2 An example ICME observed by Solar Orbiter